# Exact Exploratory Bi-factor Analysis: A Constraint-based Optimisation Approach

### v1

@@ 3.2 Study II
We conduct simulations under four settings, with $(J,G^{*})\in\{(15,3),(30,5)\}$ and $N\in\{500,2000\}$. The same data generation model in Study I is used. Under each setting, 100 independent simulations are performed. The results are given in Table 3, where the column indicated by $\bar{G}$ reports the average value of $\widehat{G}$. We see that both methods can select the number of factors reasonably well, with their accuracy being 100% when $G^{*}=3$ for both sample sizes and when $G^{*}=5$ for the larger sample size $N=2000$. This is not surprising, given that the BIC has asymptotic consistency in selecting the number of factors under both models. When $G^{*}=5$ and for the smaller sample size $N=500$, which is the most challenging setting, the proposed method achieves an accuracy of 93%, while that of the exploratory factor analysis method is only 81%. It suggests that the proposed method has a better performance in smaller sample settings. This result is expected, as the exploratory factor analysis method doesn’t utilize the information about the bi-factor structure. Consequently, it overestimates the number of parameters, which leads to a larger penalty term and, subsequently, a tendency to under-select $G$.
Table 3: Simulation results of the selection of number of factors by BIC.
@@ 4 Real Data Analysis
In this section, we apply the exact exploratory bi-factor analysis to a personality assessment dataset based on the International Personality Item Pool (IPIP) NEO 120 personality inventory (Johnson,, 2014) [^1] The data are downloaded from https://osf.io/tbmh5/. We investigate the structure of the Extraversion scale based on a sample of 1,107 UK male participants aged between 25 and 30 years. This scale consists of 24 items, which are designed to measure six facets of Extraversion, including Friendliness (E1), Gregariousness (E2), Assertiveness (E3), Activity Level(E4), Excitement-Seeking (E5) and Cheerfulness (E6); see Section B for the details. All the items are on a 1-5 Likert scale, and the reversely worded items have been reversely scored so that a larger score always means a higher level of extraversion. There is no missing data. Detailed descriptions of the items can be found in the Appendix B.
Using a candidate set $\mathcal{G}=\{2,\ldots,12\}$, the BIC procedure given in Section 2.3 selects seven group factors. The estimated loading matrix is given in Table 4, and the estimated factor correlation matrix is given below. We point out that the estimated bi-factor structure does not meet Condition 3, one of the sufficient conditions for Theorem 1. However, as shown in the Appendix C.2, with an additional mild assumption, this structure and its parameters are still identifiable.
We now examine the estimated model. We first notice that the loadings on the general factor are all positive. Consequently, this factor can be easily interpreted as the general extraversion factor. The seven group factors are closely related to the six aspects of extraversion. Specifically, we interpret the group factors G1, G3, G4 and G5 as the Friendliness, Cheerfulnes, Assertiveness, and Excitement-Seeking factors, respectively, as the items loading on them highly overlap with the items that are used to define the corresponding aspects. In particular, the items loading on G3 and G5 are exactly those that define the Cheerfulness and Excitement-Seeking aspects, respectively. The items loading on G1 include all the items that define the Friendliness aspect and an additional item “7. Prefer to be alone”, a negatively worded item that is used to define the Gregariousness aspect. This additional item aligns well with the Friendliness dimension, given the social nature behind it. In addition, the items loading on G4 consist of all the items that define the Assertiveness aspect and an additional item “6. Talk to a lot of different people at parties”, which is used to define the Gregariousness aspect. This additional item aligns with the Assertiveness dimension in that talking to many different people at parties typically requires sufficient confidence, a key element of Assertiveness.
@@
$\widehat{\Phi}=\left(\begin{matrix}1&0&0&0&0&0&0&0\\ 0&1&-0.24&0.54&0.37&0.16&0.51&0.08\\ 0&-0.24&1&-0.04&0.05&-0.09&-0.01&0.51\\ 0&0.55&-0.04&1&0.30&0.28&0.20&0.25\\ 0&0.37&0.05&0.30&1&0.38&0.15&0.29\\ 0&0.16&-0.09&0.28&0.38&1&0.11&0.22\\ 0&0.51&-0.01&0.20&0.15&0.11&1&0.10\\ 0&0.08&0.51&0.25&0.29&0.22&0.10&1\\ \end{matrix}\right).$ (10)
@@ 5 Discussions
This paper proposes a constraint-based optimisation method for exploratory bi-factor analysis. This method turns the problem of exploratory bi-factor analysis into an equality-constrained optimisation problem in a continuous domain and solves this optimisation problem by an augmented Lagrangian method. Compared with the rotation method of Jennrich and Bentler, (2011, 2012), the proposed method can learn an exact loading structure without a post-hoc treatment step. In the simulation studies, the ALM method achieves higher estimation accuracy and has a greater likelihood of recovering the entire bi-factor structure compared to the previous rotation method. Additionally, the ALM method correctly estimates the number of the group factors in most of the simulation replications. In the real data analysis concerning the Extraversion scale, the ALM method explores a bi-factor structure with seven group factors. Moreover, the selected group factors are psychologically interpretable, demonstrating the utility of our proposed method.
@@
Finally, as we demonstrate in Appendix A, the set of constraints in (1) can be extended to characterise the loading structure of a hierarchical factor model (Schmid and Leiman,, 1957; Yung et al.,, 1999), which can be used to learn a hierarchical factor structure. This exploratory hierarchical factor analysis may allow researchers to learn more refined and interpretable latent structures from psychometric data. However, one should note that exploratory hierarchical factor analysis is more complex than exploratory bi-factor analysis, as the factor hierarchy in the former can be much more complex than the two-layer hierarchy in the latter. The learning algorithm in Appendix A requires the factor hierarchy to be known (see, e.g., Panel (b) of Figure A.1). When it is unknown, one may learn the hierarchical factor structure, including the factor hierarchy as a part, by combining an iterative procedure that adaptively adds hierarchical factors into the model and a true pruning method for removing the redundant factors. Adding hierarchical factors can be done by using the learning algorithm discussed in Appendix A, and tree pruning may be achieved by a model selection procedure based on the BIC. We leave this problem for future investigation.
@@
Appendix A Extension to Hierarchical Factor Analysis
A.1 Constrained Optimisation for Exploratory Hierarchical Factor Analysis

### v2

@@ 3.2 Study II
We conduct simulations under four settings, with $(J,G^{*})\in\{(15,3),(30,5)\}$ and $N\in\{500,2000\}$ and the data generation models being the same exact bi-factor models in Study I. For each setting, 100 independent simulations are performed. The results are given in Table 4, where the column indicated by $\bar{G}$ reports the average value of $\widehat{G}$. We see that both methods can select the number of factors reasonably well, with their accuracy being 100% when $G^{*}=3$ for both sample sizes. When $G^{*}=5$ and the sample size $N=2000$, the proposed method achieves an accuracy of 100%, and the exploratory factor analysis method achieves an accuracy of 99%. This is not surprising, given that the BIC has asymptotic consistency in selecting the number of factors under both models. It is worth noting that, when $G^{*}=5$ and for the smaller sample size $N=500$, which is the most challenging setting, the proposed method achieves an accuracy of 98%, while that of the exploratory factor analysis method is zero. More precisely, the exploratory factor analysis method selects $G=4$ in all the replications. It suggests that the proposed method has an advantage in smaller sample settings. This result is expected, as the exploratory factor analysis method doesn’t utilize the information about the bi-factor structure. Consequently, it overestimates the number of parameters, which leads to a larger penalty term and, subsequently, a tendency to under-select $G$.
Table 4: Simulation results of the selection of the number of factors by BIC.
@@ 4 Real Data Analysis
In this section, we apply the exact exploratory bi-factor analysis to a personality assessment dataset based on the International Personality Item Pool (IPIP) NEO 120 personality inventory (Johnson,, 2014) [^1] The data are downloaded from https://osf.io/tbmh5/. We investigate the structure of the Extraversion scale based on a sample of 1,107 UK male participants aged between 25 and 30 years. This scale consists of 24 items, which are designed to measure six facets of Extraversion, including Friendliness (E1), Gregariousness (E2), Assertiveness (E3), Activity Level(E4), Excitement-Seeking (E5) and Cheerfulness (E6); see Section E for the details. All the items are on a 1-5 Likert scale, and the reversely worded items have been reversely scored so that a larger score always means a higher level of extraversion. There is no missing data. Detailed descriptions of the items can be found in the Appendix E.
Using a candidate set $\mathcal{G}=\{2,\ldots,12\}$, the BIC procedure given in Section 2.3 selects seven group factors. The estimated loading matrix is given in Table 5, and the estimated factor correlation matrix is given below. The estimated model fits the data well, as implied by the commonly used fit statistics, including RMSEA = 0.044, SRMR = 0.031, CFI = 0.965, and TLI = 0.953. We point out that the estimated bi-factor structure does not meet Condition 3, one of the sufficient conditions for Theorem 1. However, as shown in Appendix G.2, with some additional mild assumptions, this structure and its parameters are still identifiable.
We now examine the estimated model. We first notice that the loadings on the general factor are all positive. Consequently, this factor can be easily interpreted as the general extraversion factor. The seven group factors are closely related to the six aspects of extraversion. Specifically, we interpret the group factors G1, G3, G4 and G5 as the Friendliness, Cheerfulnes, Assertiveness, and Activity Level factors, respectively, as the items loading on them highly overlap with the items that are used to define the corresponding aspects. In particular, the items loading on G3 and G5 are exactly those that define the Cheerfulness and Activity Level aspects, respectively. The items loading on G1 include all the items that define the Friendliness aspect and an additional item “7. Prefer to be alone”, a negatively worded item that is used to define the Gregariousness aspect. This additional item aligns well with the Friendliness dimension, given the social nature behind it. In addition, the items loading on G4 consist of all the items that define the Assertiveness aspect and an additional item “6. Talk to a lot of different people at parties”, which is used to define the Gregariousness aspect. This additional item aligns with the Assertiveness dimension in that talking to many different people at parties typically requires sufficient confidence, a key element of Assertiveness.
@@
$\widehat{\Phi}=\left(\begin{matrix}1&0&0&0&0&0&0&0\\ 0&1&-0.24&0.54&0.37&0.16&0.51&0.08\\ 0&-0.24&1&-0.04&0.05&-0.09&-0.01&0.51\\ 0&0.54&-0.04&1&0.30&0.28&0.20&0.25\\ 0&0.37&0.05&0.30&1&0.38&0.15&0.29\\ 0&0.16&-0.09&0.28&0.38&1&0.11&0.22\\ 0&0.51&-0.01&0.20&0.15&0.11&1&0.10\\ 0&0.08&0.51&0.25&0.29&0.22&0.10&1\\ \end{matrix}\right).$ (11)
@@
We have also applied the bi-factor rotation method of Jennrich and Bentler, (2012) to the same data, which gives a solution with seven group factors. The resulting bi-factor structure is similar to that given by the proposed method, except that the rotation solution does not seem to contain a clear Friendliness factor; see Appendix F for further details.
@@ 5 Discussions
This paper proposes a constraint-based optimisation method for exploratory bi-factor analysis. This method turns the problem of exploratory bi-factor analysis into an equality-constrained optimisation problem in a continuous domain and solves this optimisation problem by an augmented Lagrangian method. Compared with the rotation method of Jennrich and Bentler, (2011, 2012), the proposed method can learn an exact loading structure without a post-hoc treatment step. In the simulation studies, the ALM method achieves higher estimation accuracy when data are generated from an exact bi-factor model. In addition, it has a higher chance of recovering the true bi-factor structure than the rotation method, whether data are generated from an exact or approximate bi-factor model. Moreover, the ALM method correctly estimates the number of the group factors in most of the simulation replications. In the real data analysis concerning an Extraversion personality scale, the ALM method yields a bi-factor structure with seven group factors. The identified group factors are psychologically interpretable.
@@
Third, as we demonstrate in Appendix D, the set of constraints in (1) can be extended to characterise the loading structure of a hierarchical factor model (Schmid and Leiman,, 1957; Yung et al.,, 1999), which can be used to learn a hierarchical factor structure. This exploratory hierarchical factor analysis may allow researchers to learn more refined and interpretable latent structures from psychometric data. However, one should note that exploratory hierarchical factor analysis is more complex than exploratory bi-factor analysis, as the factor hierarchy in the former can be much more complex than the two-layer hierarchy in the latter. The learning algorithm in Appendix D requires the factor hierarchy to be known (see, e.g., Panel (b) of Figure D.1). The problem becomes more challenging when the factor hierarchy is unknown, in which case we need to learn both the factor hierarchy and the loading pattern of the variables. We leave this problem for future investigation.
Finally, we point out that the proposed method always returns an estimated bi-factor model, whether it fits the data or not. The simulation study in Section 3 shows that the proposed method has robust performance when data are generated by an approximate bi-factor model. However, under more general settings, it remains to test the goodness-of-fit of the estimated model to decide whether a bi-factor model suits the data. If the bi-factor model does not fit the data well, we may consider a more flexible factor model. For example, we may apply the bi-factor rotation method or a rotation method for traditional exploratory factor analysis to allow for more cross-loadings. Alternatively, we may learn approximate bi-factor models in an exploratory manner by relaxing the equality constraints in (2) with inequality constraints in the form of $|\lambda_{jk}\lambda_{jk^{\prime}}|\leq\epsilon$, for all $j=1,...,J$, and $k,k^{\prime}=2,...,G+1$, $k\neq k^{\prime}$, where $\epsilon$ is a tuning parameter that controls the level of approximation to a bi-factor model. A larger value of $\epsilon$ leads to a more flexible model space and, thus, a more satisfactory fit, while a smaller value of $\epsilon$ leads to a better approximation to a bi-factor model that may have better interpretability. In this sense, $\epsilon$ provides a trade-off between model goodness-of-fit and bi-factor interoperability. This inequality-constrained optimisation may be solved using an interior-point method, which can incorporate the inequality constraints through suitable barrier functions (e.g., log-barriers). We leave this idea for future investigation.
@@
Appendix A Reparameterization of $\Phi$
To deal with the constraints in $\Phi$, we consider the following reparameterisation that has been considered in Alfonzetti et al., (2024), which is also similar to the implementation in the state-of-the-art statistical software Stan (Stan Development Team,, 2022):
$\Phi=\left(\begin{bmatrix}1&\mathbf{0}^{\top}\\ \mathbf{0}&U^{T}\end{bmatrix}\right)\left(\begin{bmatrix}1&\mathbf{0}^{\top}\\ \mathbf{0}&U\end{bmatrix}\right),$ (A.1)
where $U$ is defined recursively by
[…]
Appendix B Population Parameter Values in Simulations
In this section, we supplement the population values of factor loadings and factor correlations in Section 3. Under the setting $(J,G)=(15,3)$, the loading matrix $\Lambda^{*}$ and $\Phi^{*}$ in (9) and (10) are given in (B.1), (B.2), (B.3) and (B.4) respectively.
$\Lambda^{*}$ $=\left(\begin{matrix}0.36&-0.60&0&0\\ 0.89&0&0.64&0\\ 0.94&0&0&0.92\\ 0.90&0.33&0&0\\ 0.45&0&-0.90&0\\ 0.24&0&0&0.86\\ 0.45&-0.60&0&0\\ 0.74&0&-0.75&0\\ 0.64&0&0&-0.60\\ 0.50&-0.83&0&0\\ 0.54&0&-0.43&0\\ 0.43&0&0&0.86\\ 0.46&-0.33&0&0\\ 0.67&0&0.34&0\\ 0.90&0&0&-0.83\\ \end{matrix}\right)$ (B.1)
$\Phi^{*}$ $=\left(\begin{matrix}1&0&0&0\\ 0&1&0.29&0.11\\ 0&0.29&1&0.01\\ 0&0.11&0.01&1\\ \end{matrix}\right)$ (B.2)
$\Lambda^{*}$ $=\left(\begin{matrix}0.36&-0.6&-0.05&0.06\\ 0.89&0.04&0.64&0.09\\ 0.94&-0.09&-0.08&0.92\\ 0.90&0.33&0.02&-0.09\\ 0.45&-0.07&-0.90&-0.06\\ 0.24&-0.02&0.09&0.86\\ 0.45&-0.60&-0.09&0.02\\ 0.74&0.05&-0.75&0.09\\ 0.64&0.05&-0.02&-0.60\\ 0.50&-0.83&0.05&0.06\\ 0.54&0.02&-0.43&0.07\\ 0.43&0.03&-0.09&0.86\\ 0.46&-0.33&0.00&0.04\\ 0.67&0.05&0.34&0.04\\ 0.90&0.00&-0.06&-0.83\\ \end{matrix}\right)$ (B.3)
$\Phi^{*}$ $=\left(\begin{matrix}1&0&0&0\\ 0&1&-0.22&-0.07\\ 0&-0.22&1&0.46\\ 0&-0.07&0.46&1\\ \end{matrix}\right)$ (B.4)
Under the setting $(J,G)=(30,5)$, the loading matrix $\Lambda$ in (9) and (10) are given in (B.5) and (B.6). The correlation matrix $\Phi$ in (9) and (10) are given in (B.7) and (B.8).
$\footnotesize\Lambda^{*}=\left(\begin{array}[]{cccccc}0.40&-0.90&0&0&0&0\\ 0.53&0&-0.62&0&0&0\\ 0.21&0&0&-0.78&0&0\\ 0.69&0&0&0&-0.81&0\\ 0.57&0&0&0&0&0.90\\ 0.91&-0.93&0&0&0&0\\ 0.31&0&-0.47&0&0&0\\ 0.41&0&0&-0.79&0&0\\ 0.31&0&0&0&-0.35&0\\ 0.90&0&0&0&0&-0.59\\ 0.59&0.80&0&0&0&0\\ 0.83&0&-0.96&0&0&0\\ 0.84&0&0&-0.89&0&0\\ 0.29&0&0&0&-0.23&0\\ 0.90&0&0&0&0&0.77\\ 0.68&-0.56&0&0&0&0\\ 0.95&0&0.89&0&0&0\\ 0.87&0&0&0.61&0&0\\ 0.45&0&0&0&0.41&0\\ 0.52&0&0&0&0&0.51\\ 0.43&-0.42&0&0&0&0\\ 0.43&0&-0.30&0&0&0\\ 0.34&0&0&-0.31&0&0\\ 0.73&0&0&0&0.53&0\\ 0.69&0&0&0&0&-0.83\\ 0.30&0.34&0&0&0&0\\ 0.25&0&-0.60&0&0&0\\ 0.58&0&0&0.37&0&0\\ 0.28&0&0&0&-0.28&0\\ 0.85&0&0&0&0&0.82\\ \end{array}\right)$ (B.5)
$\small\Lambda^{*}=\left(\begin{array}[]{cccccc}0.40&-0.90&-0.07&0.01&0.01&0.02\\ 0.53&0.02&-0.62&-0.06&0.09&0.00\\ 0.21&0.01&-0.01&-0.78&-0.02&-0.03\\ 0.69&-0.05&0.01&0.08&-0.81&-0.07\\ 0.57&0.05&0.06&-0.09&0.03&0.90\\ 0.91&-0.93&-0.09&-0.01&0.05&0.04\\ 0.31&0.02&-0.47&0.04&0.05&-0.05\\ 0.41&0.04&-0.07&-0.79&0.02&0.05\\ 0.31&0.05&0.01&-0.07&-0.35&-0.04\\ 0.90&0.05&-0.07&-0.10&0.03&-0.59\\ 0.59&0.80&-0.05&-0.10&-0.05&0.02\\ 0.83&0.10&-0.96&-0.07&0.06&-0.03\\ 0.84&0.03&0.08&-0.89&-0.01&-0.04\\ 0.29&0.09&-0.05&-0.01&-0.23&0.04\\ 0.90&-0.10&-0.04&0.07&-0.08&0.77\\ 0.68&-0.56&0.02&-0.01&-0.07&-0.08\\ 0.95&0.05&0.89&-0.03&0.04&-0.08\\ 0.87&0.09&0.01&0.61&-0.07&0.05\\ 0.45&0.04&-0.04&0.00&0.41&0.06\\ 0.52&-0.04&-0.05&-0.10&-0.03&0.51\\ 0.43&-0.42&-0.07&-0.04&-0.07&-0.08\\ 0.43&0.02&-0.30&0.05&-0.09&0.06\\ 0.34&0.05&-0.02&-0.31&0.05&-0.04\\ 0.73&-0.07&0.10&0.09&0.53&0.04\\ 0.69&-0.09&-0.03&-0.08&-0.05&-0.83\\ 0.30&0.34&0.09&-0.04&0.01&0.01\\ 0.25&0.02&-0.60&0.09&0.04&0.06\\ 0.58&-0.08&-0.03&0.37&-0.02&0.08\\ 0.28&-0.08&0.04&-0.03&-0.28&-0.01\\ 0.85&-0.06&0.07&0.08&-0.09&0.82\\ \end{array}\right)$ (B.6)
$\Phi^{*}$ $=\left(\begin{matrix}1&0&0&0&0&0\\ 0&1&-0.37&-0.26&-0.41&-0.07\\ 0&-0.37&1&-0.08&-0.14&-0.26\\ 0&-0.26&-0.08&1&0.42&0.06\\ 0&-0.41&-0.14&0.42&1&0.20\\ 0&-0.07&-0.26&0.06&0.20&1\\ \end{matrix}\right)$ (B.7)
$\Phi^{*}$ $=\left(\begin{matrix}1&0&0&0&0&0\\ 0&1&-0.15&0.15&0.10&-0.19\\ 0&-0.15&1&0.22&-0.01&-0.03\\ 0&0.15&0.22&1&-0.03&-0.15\\ 0&0.10&-0.01&-0.03&1&0.28\\ 0&-0.19&-0.03&-0.15&0.28&1\\ \end{matrix}\right)$ (B.8)
Appendix C Sensitivity Analysis
In this section, we carry out a sensitivity analysis on the parameters $c_{\theta}$ and $c_{\sigma}$ of the proposed ALM method. We consider the same exact bi-factor model settings as in Study I of Section 3.1. For each settings, we choose $c_{\theta}\in\{0.25,0.5,0.75\}$ and $c_{\sigma}\in\{5,10,15\}$, resulting in 9 possible combinations of $(c_{\theta},c_{\sigma})$. The estimation of loading matrix, the computation time of ALM, and the results of the recovery of the bi-factor structure are shown in Table 1(d) to Table 4(d). From the sensitivity analysis, we can see that the ALM’s results are relatively stable with respect to the choice of parameters $c_{\theta}$ and $c_{\sigma}$.
Table C.1: Sensitivity Analysis of MSE of $\widehat{\Lambda}$.
$c_{\sigma}$ $c_{\theta}$ 0.25 0.5 0.75
5 $2.10\times 10^{-3}$ $2.10\times 10^{-3}$ $2.10\times 10^{-3}$
10 $2.10\times 10^{-3}$ $2.10\times 10^{-3}$ $2.10\times 10^{-3}$
15 $2.10\times 10^{-3}$ $2.10\times 10^{-3}$ $2.10\times 10^{-3}$
(a) $J=15$, $G=3$, $n=500$
$c_{\sigma}$ $c_{\theta}$ 0.25 0.5 0.75
5 $0.54\times 10^{-3}$ $0.54\times 10^{-3}$ $0.54\times 10^{-3}$
10 $0.54\times 10^{-3}$ $0.54\times 10^{-3}$ $0.54\times 10^{-3}$
15 $0.54\times 10^{-3}$ $0.54\times 10^{-3}$ $0.54\times 10^{-3}$
(b) $J=15$, $G=3$, $n=2000$
$c_{\sigma}$ $c_{\theta}$ 0.25 0.5 0.75
5 $1.36\times 10^{-3}$ $1.38\times 10^{-3}$ $1.39\times 10^{-3}$
10 $1.36\times 10^{-3}$ $1.42\times 10^{-3}$ $1.33\times 10^{-3}$
15 $1.42\times 10^{-3}$ $1.36\times 10^{-3}$ $1.44\times 10^{-3}$
(c) $J=30$, $G=5$, $n=500$
$c_{\sigma}$ $c_{\theta}$ 0.25 0.5 0.75
5 $0.30\times 10^{-3}$ $0.30\times 10^{-3}$ $0.30\times 10^{-3}$
10 $0.30\times 10^{-3}$ $0.30\times 10^{-3}$ $0.30\times 10^{-3}$
15 $0.30\times 10^{-3}$ $0.30\times 10^{-3}$ $0.30\times 10^{-3}$
(d) $J=30$, $G=5$, $n=2000$
Table C.2: Sensitivity Analysis of EMC.
$c_{\sigma}$ $c_{\theta}$ 0.25 0.5 0.75
5 1.00 1.00 1.00
10 1.00 1.00 1.00
15 1.00 1.00 1.00
(a) $J=15$, $G=3$, $n=500$
$c_{\sigma}$ $c_{\theta}$ 0.25 0.5 0.75
5 1.00 1.00 1.00
10 1.00 1.00 1.00
15 1.00 1.00 1.00
(b) $J=15$, $G=3$, $n=2000$
$c_{\sigma}$ $c_{\theta}$ 0.25 0.5 0.75
5 0.86 0.86 0.85
10 0.85 0.84 0.85
15 0.83 0.86 0.83
(c) $J=30$, $G=5$, $n=500$
$c_{\sigma}$ $c_{\theta}$ 0.25 0.5 0.75
5 1.00 1.00 1.00
10 1.00 1.00 1.00
15 1.00 1.00 1.00
(d) $J=30$, $G=5$, $n=2000$
Table C.3: Sensitivity Analysis of ACC.
$c_{\sigma}$ $c_{\theta}$ 0.25 0.5 0.75
5 1.000 1.000 1.000
10 1.000 1.000 1.000
15 1.000 1.000 1.000
(a) $J=15$, $G=3$, $n=500$
$c_{\sigma}$ $c_{\theta}$ 0.25 0.5 0.75
5 1.000 1.000 1.000
10 1.000 1.000 1.000
15 1.000 1.000 1.000
(b) $J=15$, $G=3$, $n=2000$
$c_{\sigma}$ $c_{\theta}$ 0.25 0.5 0.75
5 0.998 0.998 0.998
10 0.998 0.997 0.998
15 0.997 0.998 0.997
(c) $J=30$, $G=5$, $n=500$
$c_{\sigma}$ $c_{\theta}$ 0.25 0.5 0.75
5 1.000 1.000 1.000
10 1.000 1.000 1.000
15 1.000 1.000 1.000
(d) $J=30$, $G=5$, $n=2000$
Table C.4: Sensitivity Analysis of Computation time(s).
$c_{\sigma}$ $c_{\theta}$ 0.25 0.5 0.75
5 0.13 0.13 0.13
10 0.13 0.13 0.12
15 0.10 0.09 0.08
(a) $J=15$, $G=3$, $n=500$
$c_{\sigma}$ $c_{\theta}$ 0.25 0.5 0.75
5 0.10 0.09 0.10
10 0.10 0.09 0.09
15 0.07 0.06 0.06
(b) $J=15$, $G=3$, $n=2000$
$c_{\sigma}$ $c_{\theta}$ 0.25 0.5 0.75
5 0.52 0.51 0.50
10 0.49 0.47 0.41
15 0.36 0.33 0.30
(c) $J=30$, $G=5$, $n=500$
$c_{\sigma}$ $c_{\theta}$ 0.25 0.5 0.75
5 0.38 0.37 0.37
10 0.37 0.33 0.28
15 0.23 0.22 0.24
(d) $J=30$, $G=5$, $n=2000$
Appendix D Extension to Hierarchical Factor Analysis
D.1 Constrained Optimisation for Exploratory Hierarchical Factor Analysis